OpenStudy (anonymous):

Solve 2|2x + 2| + 7 = 35 for x. Show your work to receive full credit.

OpenStudy (anonymous):

\[ 2|2x + 2| + 7 = 35\\ 2|2x + 2| = 28\\ 4|x +1| = 28\\ |x +1| = 7\\ x+1 =\pm 7\\ x =6 \text { or } x=-8 \]